# Learning in Compact Spaces with Approximately Normalized Transformers

### v1

@@ 4.3 Approximal normalized GPT (anGPT)
In the following, we describe our approximal normalized GPT model, anGPT inspired by [5]. In contrast to nGPT, we perform normalization consistent with the assumption above, while nGPT normalizes along the residual dimension (to stay on a hypersphere). This leads to different normalization dimensions for linear maps in the input and output of an attention or feed-forward layer. We define $d_{m}$, $d_{h}$, $d_{f}$ as the model, attention head, and up-scaled dimension respectively. In the feed-forward MLP, $l$ as the number of layers, and $d_{v}$ as the vocabulary size. An overview of our architectural modifications, including nGPT, can be found in Table 1.
@@
First of all, we remove all classical normalization layers and replace them with a post L2 normalization $\text{norm}(x)=x/\lVert x\rVert_{2}$. Removing normalization completely leads to unstable training. We replace the classic residual update by LERP and replace the learnable per-element affine parameter from the pre-norm layer by a learnable interpolation parameter $\bm{\alpha}$.
@@
We add constant normalization factors in the attention layer for the query, key, and value map, but due to the reshape into the head dimension, we consider the head dimension as the output dimension $\nu_{qkv}=\sqrt{d_{m}/d_{h}}$. Similarly, we add a normalization factor for the output map $\nu_{p}=\sqrt{d_{h}/d_{m}}$ but with $d_{h}$ as input dimension. In the feed forward layer we add three normalization factors $\nu_{uz}=\sqrt{d_{m}/(4d_{m})}$, $\nu_{d}=\sqrt{(4d_{m})/d_{m}}$, and $\approx 3.74$ (estimated via Monte Carlo) for upscaling, downscaling and the activation function, respectively. The normalization factor for the residual LERP update is a function of $\bm{\alpha}$ and calculated at each step through $\nu(\bm{\alpha})=1-2\bm{\alpha}+2\bm{\alpha}^{2}$. During inference, we can subsume constant normalization factors and the logits scaling into the model parameters. Since we need no more normalization layers than the GPT model, we expect the same inference time.
@@
In anGPT, we remove the RMSnorm before the head linear since the representation is already normalized. However, to scale the logits, we add a learnable scaling vector $\bm{s_{z}}\in\mathbb{R}^{d_{v}}$ similar to nGPT. We also tried to scale before the head linear, but it reduced performance.
Scale effective learning rates
Similar to Loshchilov et al. [5], we increase the effective learning rate of interpolation vector $\bm{\alpha}$ and scaling parameters $s_{z}$ by learning a scale instead of the parameter directly. Therefore, we use two scalars $s_{init}$ and $s_{scale}$. We initialize the learnable parameter with $s_{scale}$ and multiply the parameter in each forward pass with $s_{init}/s_{scale}$. Setting the initial parameter scale to $s_{init}=1/\sqrt{d_{m}}$ leads to an effective learning rate similar to weight parameters in the network. However, we found that further increasing the effective learning rate benefits the training progress.
Figure 2: Scaling trend fits for optimal batch size and learning rate as functions of model size $N$. Grid point markers are shaded by excess loss relative to all configurations for this parameter. Diamond markers show the two-stage interpolation-based estimates of optimal hyperparameters. Dashed lines represent fitted power laws using the estimated optimal hyperparameters.
@@ 5 Experiments
In the following, we describe multiple LLM pretraining experiments comparing anGPT to GPT and nGPT. We use SlimPajama [14] with ~$\text{\textasciitilde}627B$ token and train for $<1$ epoch. For all experiments, we use the GPT-NeoX tokenizer with a 50k vocabulary size [15] and a context window of 2048 tokens. As a baseline, we extend the vanilla GPT architecture with a SwiGLU activation function [16], rotary position embedding [17], RMS normalization [8], and QK normalization [18] and dub the resulting model GPT+. We adjust the batch size according to the model size to get an optimal GPU utilization. We train all models with Adam [11] and, in the case of GPT+, with AdamW [19]. For learning rate scheduling, we use a cosine learning rate annealing [20]. If learning rate warm-up is used, we dedicate $10\%$ of the total training steps as warm-up. In the following experiments, we compare anGPT to nGPT and GPT+. We report hyperparameters in Appendix F and comparison to related work in Appendix A.
@@ 5.1 Hyperparameter Scaling Trends
To use the right learning rate and batch size for the baseline and our approach, we derived hyperparameter scaling trends across model sizes from $N=32M$ to $N=0.5B$ parameters (Configuration in Table F.1). We orient our methodology on Porian et al. [21] and DeepSeek-AI et al. [10]. We performed a grid search for each model size over the batch size and learning rate with a Chinchilla optimal token budget ($D=20N$, [22]). Optimal batch size and learning rate for each model size were obtained via a two-stage interpolation process adapted from Porian et al. [21]. The estimates for the optimal values are shown in Figure 2. We describe the process in detail in Appendix G and provide the full hyperparameter sweep results in Figure G.1.
When fitting hyperparameter scaling trends for the optimal parameter selection, we found that anGPT can use larger batch sizes for larger model sizes, which could be beneficial for scaling the pretraining to a large number of workers for achieving speedup. For the scaling behavior of the learning rate, we found a small negative exponent for anGPT, so the learning rate decreases gradually as models get larger, which is expected behavior in line with previous works [23, 10].
Figure 3: Training the $0.5B$ model up to 7×7$\times$7 × Chinchilla optimal token budget. Each point is the final validation loss of a full training run with the training budget noted on the abscissa. Below, we measure the convergence speed-up against the GPT+ model with QK normalization.
@@ 5.2 Performance comparison
To compare our approach to GPT+ and nGPT, we trained a $0.5B$ model of each architecture on token budgets from $5B$ to $70B$ (0.5×0.5$\times$0.5 × to 7×7$\times$7 × Chinchilla optimal). For GPT+ and anGPT, we used the optimal hyperparameters from the previous grid search and performed an additional grid search for nGPT. We report the results and the convergence speedup against GPT+ in Figure 3. We found that anGPT and nGPT outperform GPT+ on any token budget, while for smaller token budgets, nGPT performs better and reaches on par performance for larger ones. When measuring the convergence improvement to GPT+, we found an average speedup factor of 1.4×1.4$\times$1.4 × for anGPT and 1.29×1.29$\times$1.29 × for nGPT. Similar to Loshchilov et al. [5], we measured the convergence speedup against GPT+ without QK normalization and achieved speedup factors of 2.0×2.0$\times$2.0 × for anGPT and 1.8×1.8$\times$1.8 × for nGPT, see Figure H.2. We further discuss the difference to their reported speedups in Appendix H.
In Table 2, we report the average runtime per training step for all three architectures. We found an increase for anGPT of ~$\text{\textasciitilde}3\%$ and ~$\text{\textasciitilde}9\%$ for nGPT. We attribute the additional runtime to the additional scaling factors, the norm implementation, and the additional norm for nGPT. When running inference, we found similar run times for GPT+ and anGPT but a ~$\text{\textasciitilde}3\%$ increase for nGPT, due to the additional normalization operations.
Table 2: Runtime Comparison with a $0.5B$ parameter model on a GPU node with 4 A100 (40GB) GPUs with a sequence length of 2048 and a batch size of 8. We use Flash Attention $0.7.4$ and PyTorch $2.6$ with torch.compile default settings.
@@ 5.3 Comparing performance via compute-dependent scaling law
We further compare GPT+ and anGPT by training model sizes from $32M$ to $1B$ parameters on token budgets from 0.5×0.5$\times$0.5 × to 5×5$\times$5 × Chinchilla optimal. This allows us to investigate the scaling behavior of the new anGPT architecture. We report the results and scaling fits in Figure 4. Given the same compute budget, anGPT outperforms GPT+ on any model size and compute budget. We further used the results to derive scaling laws using the approach of Hoffmann et al. [22] and described in Appendix I. We found that both GPT+ and anGPT have nearly identical estimated scaling law exponents, which implies that the improvement of validation loss is not significantly different across these architectures.
Figure 4: Training different model sizes on different token budgets. Each point is a full training with the training budget noted on the abscissa. We fit for both architectures a scaling law as described in Appendix I and indicated by the dashed line.
5.4 Ablation study
To investigate the modification we add to a vanilla GPT architecture, we run small experiments on a $0.5B$ GPT model trained on $10B$ tokens from OpenWebText [24]. When adding QK norm to the baseline, it leads to a performance gain of $2.1\%$ as visualized in Figure 5. We get an additional benefit by using nGPT and can further improve it by using weight bounds instead of weight normalizations. The replacement of the normalization layer after the LERP update by a normalization factor does not reduce the performance. For anGPT (additional normalization factors, different normalization dimension, removing scaling vectors), we see an additional gain of $0.4\%$.
@@ 5.4 Ablation study
Figure 5: We run ablation experiments with a $0.5B$ parameter model using $10B$ tokens from OpenWebText. We see a gain by adding QK norm. We modify nGPT by replacing the normalization of the LERP update with a normalization factor and, in addition, by bounding weights instead of normalizing them. The anGPT mainly replaces scaling vectors by normalization factors.
@@ 6 Limitations
Despite the extensive experimental evaluation, we do not perform experiments with more than 7×7$\times$7 × Chinchilla optimal token budgets; our approach could perform worse than nGPT or GPT+ training with larger budgets. We also did not perform multiple experiments with different random seeds to generate error bars due to the high cost of GPT pretraining. We do not evaluate on the downstream task and assume that the validation loss correlates with the downstream performance. It is also hard to get a meaningful downstream signal from small and short-trained LLMs. Lastly, we perform the majority of the experiments on only one dataset and only with the GPT architecture; the performance could diverge with different datasets or data modalities.
7 Discussion & Conclusion
In our preliminaries, we hypothesized that the benefits of normalization stem from two effects: stabilizing input scales across layers and preventing representation norm escalation. Our experimental results provide strong evidence supporting both hypotheses. We theorized that normalization improves optimization, which is supported by Figure B.1 showing that GPT exhibits significantly reduced variance in Adam’s first moments compared to GPT+, without requiring learning rate warm-up. Our second hypothesis addressed the "Curse of Depth". As seen in Figure 1, GPT+ shows input norms growing exponentially with depth, while anGPT constrains this growth. We performed extensive experiments demonstrating the strengths of anTransformer and investigating its scaling behavior.
@@ 7 Discussion & Conclusion
Our experiments confirm that anGPT effectively captures normalization benefits with minimal computational overhead. The key insight of our approach is decoupling stabilization from loss minimization. Traditional normalization layers employ learnable parameters $\gamma$ that serve dual purposes: stabilizing network activations and simultaneously participating in loss minimization. This conflation of roles complicates optimization dynamics. In anGPT, we deliberately decouple these concerns: normalization factors $\nu$ handle stabilization exclusively, while the other parameters focus on minimizing the loss. This decoupling eliminates the need for warm-up while preserving favorable scaling properties. Since we also do not require weight decay, the practical impact is substantial: hyperparameter tuning becomes simpler and scaling law derivation more efficient, as fewer parameters require optimization at each measurement point.
Due to the consistent application of normalizations and our weight matrices being bounded along the input dimensions (see Figure E.1), all representations are contained in a compact space. Our experiments show that anGPT achieves 1.4×1.4$\times$1.4 × faster convergence compared to GPT with QK norm with minimal computational overhead, see Figure 3. Overall, anTransformer demonstrates that the benefits of consistent normalization, such as a convergence speedup and fewer hyperparameters, can be achieved with minimal computational overhead, and provides a promising approach to training large language models with predictable scaling behavior.

### v2

@@ 4.3 Approximal normalized GPT (anGPT)
In the following, we describe our approximal normalized GPT model, anGPT, inspired by [loshchilov2025ngpt]. In contrast to nGPT, we perform normalization consistent with the assumption above, while nGPT normalizes along the residual dimension (to stay on a hypersphere). This leads to different normalization dimensions for linear maps in the input and output of an attention or feed-forward layer. Let $d_{m}$, $d_{h}$, $d_{f}$ denote the model, attention head, and up-scaled dimension, respectively. In the feed-forward MLP, $l$ as the number of layers, and $d_{v}$ as the vocabulary size. An overview of our architectural modifications, including nGPT, can be found in Table 1.
@@
First of all, we remove all classical normalization layers and replace them with a post-L2-normalization $\text{norm}(x)=x/\lVert x\rVert_{2}$. Removing normalization completely leads to unstable training. The architecture replaces the classic residual update with LERP and replaces the learnable per-element affine parameter from the pre-norm layer with a learnable interpolation parameter $\bm{\alpha}$.
@@
We add constant normalization factors in the attention layer for the query, key, and value map, but due to the reshape into the head dimension, we consider the head dimension as the output dimension $\nu_{qkv}=\sqrt{d_{m}/d_{h}}$. Similarly, we add a normalization factor for the output map $\nu_{p}=\sqrt{d_{h}/d_{m}}$ but with $d_{h}$ as input dimension. In the feed forward layer we add three normalization factors $\nu_{uz}=\sqrt{d_{m}/(4d_{m})}$, $\nu_{d}=\sqrt{(4d_{m})/d_{m}}$, and $\approx 3.74$ (estimated via Monte Carlo) for upscaling, downscaling and the activation function, respectively. The normalization factor for the residual LERP update is a function of $\bm{\alpha}$ and calculated at each step through $\nu(\bm{\alpha})=1-2\bm{\alpha}+2\bm{\alpha}^{2}$. A list of normalization factors for each primitive can be found in Appendix E. During inference, we can subsume constant normalization factors and the logits scaling into the model parameters. Since we need no more normalization layers than the GPT model, we expect the same inference time.
@@
anGPT removes the RMSnorm before the head linear since the representation is already normalized. However, to scale the logits, a learnable scaling vector $\bm{s_{z}}\in\mathbb{R}^{d_{v}}$ is added, similar to nGPT. Scaling before the head linear was also tested but reduced performance.
Parameter Reparameterization for Uniform Optimization
Following nGPT [loshchilov2025ngpt], we employ a reparameterization scheme to ensure uniform optimization dynamics across parameter types. For any trainable scaling parameter $s_{a}$ (e.g., $\alpha_{A}$, $\alpha_{M}$, $s_{z}$), we optimize a surrogate parameter $\hat{s}_{a}$ and compute:
$s_{a}=\frac{s_{a,\text{init}}}{s_{a,\text{scale}}}\cdot\hat{s}_{a}$
The surrogate is initialized as $\hat{s}_{a}^{(0)}=s_{a,\text{scale}}$, ensuring $s_{a}^{(0)}=s_{a,\text{init}}$. This reparameterization ensures all stored parameters have comparable magnitude $\sim s_{a,\text{scale}}$, enabling Adam’s adaptive learning rate mechanism to work uniformly across the network. The effective learning rate for updates to $s_{a}$ becomes $\alpha_{\text{eff}}=\alpha\left(s_{a,\text{init}}/s_{a,\text{scale}}\right)^{2}$.
Figure 2: Scaling trend fits for optimal batch size and learning rate as functions of model size NN. Grid point markers are shaded by excess loss relative to all configurations for this parameter. Diamond markers show the two-stage interpolation-based estimates of optimal hyperparameters. Dashed lines represent fitted power laws using the estimated optimal hyperparameters.
@@ 5 Experiments
In the following, we describe multiple LLM pretraining experiments comparing anGPT to GPT and nGPT. The experiments use SlimPajama [cerebras2023slimpajama] with $\text{\textasciitilde}627B$ tokens and train for <1<1 epoch. All experiments employ the GPT-NeoX tokenizer with a 50k vocabulary size [gpt-neox-20b] and a context window of 2048 tokens. As a baseline, we extend the vanilla GPT architecture with a SwiGLU activation function [shazeer2020glu], rotary position embedding [su2021roformer], RMS normalization [zhang2019rms], and QK normalization [henry2020qknorm] and dub the resulting model GPT+. All models are trained with Adam [kingma-iclr15a] and, in the case of GPT+, with AdamW [loshchilov-iclr19a]. For learning rate scheduling, a cosine learning rate annealing [loshchilov-iclr17a] is employed. If learning rate warm-up is used, $10\%$ of the total training steps are dedicated to warm-up. For adapting the effective learning rate of scaling parameters, we set $s_{a,\text{init}}=0.01$ for anGPT and $s_{a,\text{init}}=1/\sqrt{(}d)$ for nGPT. The following experiments compare anGPT to nGPT and GPT+. Hyperparameters are reported in Appendix F and comparisons with related work in Appendix A.
@@ 5.1 Hyperparameter Scaling Trends
To determine optimal learning rates and batch sizes, hyperparameter scaling trends were derived across model sizes from $N=32M$ to $N=0.5B$ parameters (Configuration in Table F.1). The methodology follows porian2024resolving and deepseekai2024v1. We performed a grid search for each model size over the batch size and learning rate with a Chinchilla optimal token budget ($D=20N$, [hoffmann2022compute]). Optimal batch size and learning rate for each model size were obtained via a two-stage interpolation process adapted from porian2024resolving. The estimates for the optimal values are shown in Figure 2. The process is described in detail in Appendix G with full hyperparameter sweep results in Figure G.1. When fitting hyperparameter scaling trends for the optimal parameter selection, we found that anGPT can use larger batch sizes for larger model sizes, which could be beneficial for scaling the pretraining to a large number of workers for achieving speedup. For the scaling behavior of the learning rate, we found a small negative exponent for anGPT, so the learning rate decreases gradually as models get larger, which is expected behavior in line with previous works [bai2023qwentechnicalreport, deepseekai2024v1].
Figure 3: Training the $0.5B$ model up to $7\times$ Chinchilla optimal token budget. Each point is the final validation loss of a full training run with the training budget noted on the abscissa. Below, we measure the convergence speed-up against the GPT+ model with QK normalization.
@@ 5.2 Performance comparison
To compare our approach to GPT+ and nGPT, we trained a $0.5B$ model of each architecture on token budgets from $5B$ to $70B$ ($0.5\times$ to $7\times$ Chinchilla optimal). For GPT+ and anGPT, we used the optimal hyperparameters from the previous grid search and performed an additional grid search for nGPT. Figure 3 reports the results and convergence speedup against GPT+. Results show that anGPT and nGPT outperform GPT+ across all token budgets, with nGPT performing better at smaller token budgets and reaching comparable performance at larger ones. Convergence measurements reveal an average speedup factor of $1.4\times$ for anGPT and $1.29\times$ for nGPT compared to GPT+. Similar to loshchilov2025ngpt, convergence speedup measurements against GPT+ without QK normalization yield speedup factors of $2.0\times$ for anGPT and $1.8\times$ for nGPT, see Figure H.2. We discuss the discrepancy to [loshchilov2025ngpt] reported speedups in Appendix H. In addition, we performed performance comparisons on different token budgets for the $250M$ and $1.0B$ model sizes and find the  40% convergence speedup observed for the 0.5B model consistent. Convergence plots for both models are provided in Appendix I.
Table 2 reports the average runtime per training step for all three architectures. anGPT shows an increase of $\text{\textasciitilde}3\%$ and nGPT shows $\text{\textasciitilde}9\%$ increase. The additional runtime is attributed to the additional scaling factors, the norm implementation, and the additional norm for nGPT. During inference, similar runtimes were observed for GPT+ and anGPT but a $\text{\textasciitilde}3\%$ increase for nGPT, due to the additional normalization operations.
Table 2: Runtime Comparison with a $0.5B$ parameter model on a GPU node with 4 A100 (40GB) GPUs with a sequence length of 2048 and a batch size of 8. The experiments use torch.compile with default settings.
@@ 5.3 Comparing performance via compute-dependent scaling law
We further compare GPT+ and anGPT by training model sizes from $32M$ to $1B$ parameters on token budgets from $0.5\times$ to $5\times$ Chinchilla optimal. This enables investigation of the scaling behavior of the new anGPT architecture. Figure 4 shows the results and scaling fits. Given the same compute budget, anGPT outperforms GPT+ on any model size and compute budget. These results were used to derive scaling laws using the approach of hoffmann2022compute, as described in Appendix J. Both GPT+ and anGPT exhibit nearly identical estimated scaling law exponents, implying that the improvement of validation loss is not significantly different across these architectures.
Figure 4: Training different model sizes on different token budgets. Each point represents a full training with the training budget noted on the abscissa. The scaling law is fitted for both architectures as described in Appendix J and indicated by the dashed line.
5.4 Downstream Evaluation
To verify that pretraining improvements transfer beyond perplexity metrics, we evaluate our models on standard benchmarks. The 1B anGPT model consistently outperforms GPT+ across six benchmarks and three training budgets, with improvements ranging from 3% to 22% depending on the task. Detailed downstream evaluation results are presented in Appendix K.
5.5 Ablation studies
To investigate the modifications added to a vanilla GPT architecture, small experiments were conducted on a $0.5B$ GPT model trained on $10B$ tokens from OpenWebText [Gokaslan2019OpenWeb]. When adding QK norm to the baseline, it leads to a performance gain of $2.1\%$ as visualized in Figure 5. Additional benefits emerge from using nGPT, with further improvements from using weight bounds instead of weight normalizations. The replacement of the normalization layer after the LERP update by a normalization factor does not reduce the performance. For anGPT (additional normalization factors, different normalization dimension, removing scaling vectors), we see an additional gain of $0.4\%$. Extensive ablation studies further assess the sensitivity of anGPT to variations in normalization factors. Our analysis shows that while the method is robust to small estimation errors (below 2%), the residual normalization factors are critical for training stability. Complete ablation results, including sensitivity to factor scaling and comparisons with different configurations, are provided in Appendix L.
@@ 5.5 Ablation studies
Figure 5: We run ablation experiments with a $0.5B$ parameter model using $10B$ tokens from OpenWebText. Adding QK norm shows a performance gain. We modify nGPT by replacing the normalization of the LERP update with a normalization factor and, in addition, by bounding weights instead of normalizing them. The anGPT mainly replaces scaling vectors by normalization factors.
5.6 Analysis of anGPT
[…]
@@ 6 Limitations
Despite the extensive experimental evaluation, we do not perform experiments with more than $7\times$ Chinchilla optimal token budgets; our approach could perform worse than nGPT or GPT+ training with larger budgets. We also did not perform multiple experiments with different random seeds to generate error bars due to the high cost of GPT pretraining. We do not evaluate on the downstream task and assume that the validation loss correlates with the downstream performance. It is also hard to get a meaningful downstream signal from small and short-trained LLMs. Lastly, we perform the majority of the experiments on only one dataset and only with the GPT architecture; the performance could diverge with different datasets and data modalities.
7 Discussion
@@
In our preliminaries, we hypothesized that the benefits of normalization stem from two effects: stabilizing input scales across layers and preventing representation norm escalation. Our experimental results provide strong evidence supporting both hypotheses. First, we hypothesized that normalization enhances optimization, as confirmed by Figure B.1. This shows that anGPT exhibits significantly reduced variance in Adam’s first moments compared to GPT+, without requiring learning rate warm-up. Second, our hypothesis addressed the "Curse of Depth." As seen in Figure 1, GPT+ shows input norms growing exponentially with depth, while anGPT successfully constrains this growth. Traditional normalization layers employ learnable parameters $\gamma$ that serve two purposes: stabilizing network activations while simultaneously participating in loss minimization. This conflation of roles complicates optimization dynamics. In anGPT, we deliberately decouple these concerns—normalization factors $\nu$ handle stabilization exclusively, while the remaining parameters focus solely on minimizing the loss. This decoupling eliminates the need for warm-up while preserving favorable scaling properties. Since weight decay is also unnecessary, the practical impact is substantial: hyperparameter tuning becomes simpler and scaling law derivation more efficient, as fewer parameters need to be tuned.
Interestingly, the theoretical assumptions motivating our normalization factors do not strictly hold during training and are even violated by design. This occurs because weights are bounded rather than normalized, leading to compact rather than perfectly normalized representations. Nevertheless, we still observe concentration empirically (see Figure P.1), particularly due to the high dimensionality of representations. This robustness to assumption violations actually strengthens our approach, demonstrating that exact normalization is unnecessary to achieve the benefits traditionally associated with normalization layers. The key insight is that decoupling stabilization from loss minimization enables $1.4\times$ faster convergence compared to GPT with QK norm, while adding only minor computational overhead (see Figure 3).
Looking forward, several directions warrant further investigation. The compact representation space maintained by anGPT makes it particularly amenable to reduced-precision training. The bounded nature of activations and weights could enable efficient FP8 training without the numerical instabilities typically associated with low-precision arithmetic in unbounded architectures. Additionally, extending the approximate normalization framework to other architectures such as vision transformers and diffusion models could yield similar efficiency gains. Another promising avenue is exploring whether alternative optimizers might be more suitable for normalized architectures than Adam, as the bounded parameter space and stable gradient flow could benefit from optimization algorithms specifically designed for compact spaces.
8 Conclusion
We presented anGPT, an approximately normalized transformer that achieves faster convergence through scalar multiplication-based normalization. By leveraging the concentration of norms in high-dimensional spaces and decoupling stabilization from loss minimization, our method eliminates the need for weight decay and learning rate warm-up while maintaining only 3% computational overhead. Overall, anTransformer demonstrates that the benefits of consistent normalization, such as convergence speedup and fewer hyperparameters, can be achieved with minimal computational overhead, and provides a promising approach to training large language models with predictable scaling behavior.